A 5.5 percent coupon municipal bond has 16 years left to maturity and has a price quote of 92.55. The bond can be called in nine years. The call premium is one year of coupon payments. Compute the bond's yield to maturity and yield to call. Assume interest payments are paid semiannually and a par value of $5,000. Multiple Choice A.) YTM = 6.91 percent; YTC = 7.52 percent B.) YTM = 6.24 percent; YTC = 7.08 percent C.) YTM = 5.78 percent; YTC = 6.61 percent D.) YTM = 5.92 percent; YTC = 6.85 percent
Answer #1

To find the YTM, we need to put the following values in the financial calculator:

N = 16*2 = 32;

PV = -0.9255 * 5000 = -4627.50

PMT = (5.5%/2)*5000 = 137.50

FV = 5000;

Press CPT, then I/Y, which gives us 3.12

Periodic Rate = 3.12%

So, YTM = Periodic Rate * No. of compounding periods in a year = 3.12% * 2 = 6.24%

To find the YTC, we need to put the following values in the financial calculator:

N = 9*2 = 18;

PV = -0.9255 * 5000 = -4627.50

PMT = (5.5%/2)*5000 = 137.50

FV = 5000 + (137.50*2) = 5275;

Press CPT, then I/Y, which gives us 3.54

Periodic Rate = 3.54%

So, YTC = Periodic Rate * No. of compounding periods in a year = 3.54% * 2 = 7.08%

Hence, Option "B" is correct.